# Closed-Form Interpretation of Neural Network Latent Spaces with Symbolic Gradients

### v3

@@ A.1.1 Invariants of Rank-Two Tensors Under Similarity Transformations
For any $n\times n$ matrix $A$, we can compute the similarity transform $B=MAM^{-1}$, where $M$ is any invertible matrix. Using the cyclic property of trace,
$\text{tr}(B)=\text{tr}(MAM^{-1})=\text{tr}(AM^{-1}M)=\text{tr}(A).$
@@
A.1.2 $3\times 3$ Antisymmetric Matrices
In the case of antisymmetric $n\times n$ matrices of odd size, the number of principal invariants reduces to one. The trace of an antisymmetric matrix is 00, so $I_{1}=\text{tr}(A)=0$, and any antisymmetric square matrix of odd size $n$ must have at least one zero-eigenvalue, so $I_{3}=\det(A)=0$. We treat the case of a $3\times 3$ antisymmetric matrix, in which case $I_{2}$ can be written in terms of its entries as,
@@ A.1.2 $3\times 3$ Antisymmetric Matrices
Since the diagonal elements of $A$ are 00 and $A_{ij}=-A_{ji}$, the expression for $I_{2}$ is simplified:
@@
A.1.3 Invariants of the Field Strength Tensor Under the Lorentz Transformation
Under Lorentz transformations, the invariants of the electromagnetic field strength tensor $F_{\mu\nu}$ are preserved. The tensor $F_{\mu\nu}$ is antisymmetric, meaning $F_{\mu\nu}=-F_{\nu\mu}$. Its invariants include the scalar $\mathbf{B}\cdot\mathbf{E}$ and the quantity $\frac{1}{2}F_{\mu\nu}F^{\mu\nu}$. Specifically,
@@ A.1.3 Invariants of the Field Strength Tensor Under the Lorentz Transformation
These invariants highlight the consistency of electromagnetic properties across different inertial frames.
A.2 Space-Time Interval
@@
A.3 Dynamical Systems
@@
In dynamical systems involving motion in a potential, conservation principles are fundamental. In one-dimensional systems, energy is invariant, meaning the total energy—comprising both kinetic and potential components—remains constant in an isolated system. In two-dimensional systems, both energy and angular momentum are conserved, provided the potential is central (i.e., depends only on the radial distance). These invariants are crucial for understanding and modeling dynamical behaviors in both 1D and 2D contexts.

### v4

@@ C.1.1 Invariants of Rank-Two Tensors Under Similarity Transformations
For any $n\times n$ matrix $A$, we can compute the similarity transform $B=CAC^{-1}$, where $C$ is any invertible matrix. Using the cyclic property of trace,
$\text{tr}(B)=\text{tr}(CAC^{-1})=\text{tr}(AC^{-1}M)=\text{tr}(A).$
@@
In this case, the placeholder operator $M$ is $M(A)=CAC^{-1}$.
C.1.2 $3\times 3$ Antisymmetric Matrices
In the case of antisymmetric $n\times n$ matrices of odd size, the number of principal invariants reduces to one. The trace of an antisymmetric matrix is 0, so $I_{1}=\text{tr}(A)=0$, and any antisymmetric square matrix of odd size $n$ must have at least one zero-eigenvalue, so $I_{3}=\det(A)=0$. We treat the case of a $3\times 3$ antisymmetric matrix, in which case $I_{2}$ can be written in terms of its entries as,
@@ C.1.2 $3\times 3$ Antisymmetric Matrices
Since the diagonal elements of $A$ are 0 and $A_{ij}=-A_{ji}$, the expression for $I_{2}$ is simplified:
@@
For the case of antisymmetric matrices it is important to restrict the transformation $M(A)=CAC^{-1}$ to be mediated by orthogonal matrices $C\in O(3)$ to preserve antisymmetricity.
C.1.3 Invariants of the Field Strength Tensor Under the Lorentz Transformation
The electromagnetic field strength tensor unifies the magnetic $\mathbf{B}$ and electric fields $\mathbf{E}$
$F^{\mu\nu}=\begin{bmatrix}0&-E_{x}/c&-E_{y}/c&-E_{z}/c\\ E_{x}/c&0&-B_{z}&B_{y}\\ E_{y}/c&B_{z}&0&-B_{x}\\ E_{z}/c&-B_{y}&B_{x}&0\end{bmatrix}$ (21)
where $c=1$ is the speed of light. The tensor $F_{\mu\nu}$ is antisymmetric, meaning $F_{\mu\nu}=-F_{\nu\mu}$. Under Lorentz transformations, the invariants of the electromagnetic field strength tensor $F_{\mu\nu}$ are preserved. Its invariants include the scalar $\mathbf{B}\cdot\mathbf{E}$ and the quantity $\frac{1}{2}F_{\mu\nu}F^{\mu\nu}$. Specifically,
@@ C.1.3 Invariants of the Field Strength Tensor Under the Lorentz Transformation
These invariants highlight the consistency of electromagnetic properties across different inertial frames. The placeholder operator implements $M(A)=\Lambda A\Lambda^{\top}$, where $\Lambda$ is the Lorentz transformation.
C.2 Space-Time Interval
@@
In this case, the placeholder operator performs $M(A)=\Lambda A$, where $\Lambda$ is the Lorentz transformation.
C.3 Dynamical Systems
@@
In dynamical systems involving motion in a potential, conservation principles are fundamental. In one-dimensional systems, energy is invariant, meaning the total energy—comprising both kinetic and potential components—remains constant in an isolated system. In two-dimensional systems, both energy and angular momentum are conserved, provided the potential is central (i.e., depends only on the radial distance). These invariants are crucial for understanding and modeling dynamical behaviors in both 1D and 2D contexts. The operator $M$ evolves the system by mapping a state vector $\mathbf{x}$ to its time-evolved counterpart, $M(\mathbf{x})$, representing the system’s state at a later time.
Appendix D Implementation Details
D.1 Symbolic Regression Hyperparameters
To reproduce this experiment, the SymbolicRegression.jl library was used to perform symbolic search with a custom loss function targeting gradient alignment. The model was configured with binary operators (+, -, *, /, $\hat{}$, div) and unary operators (sqrt, square, sin, exp). Complexity penalties were assigned as follows: constants had a complexity of 3, while operators had complexities of sqrt => 4, square => 4, sin => 5, and exp => 5. The training process involved niterations=200, batch_size=25, early_stop_condition=1e-10, and a maxsize=25 constraint on the equation size. Simplification of equations, optimization of constants, and automatic differentiation were enabled to improve the accuracy and interpretability of the resulting expressions.
To reproduce this experiment, the SymbolicRegression.jl library was used to perform symbolic search with a custom loss function targeting gradient alignment. The model was configured with binary operators (+, -, *, /, $\hat{}$, div) and unary operators (sqrt, square, sin, exp). Complexity penalties were assigned as follows: constants had a complexity of 3, while operators had complexities of sqrt => 4, square => 4, sin => 5, and exp => 5. The training process involved niterations=200, batch_size=25, early_stop_condition=1e-10, and a maxsize=25 constraint on the equation size. Simplification of equations, optimization of constants, and automatic differentiation were enabled to improve the accuracy and interpretability of the resulting expressions.
D.2 Training Hyperparameters
All experiments use the Adam optimizer and the scheduler class ReduceLROnPlateau from the PyTorch library. For all experiments, the sub-network $f$ is a fully-connected feedforward network designed as follows:
  • An input layer
  • Two hidden layers
  • An output layer with a single neuron
The layer sizes for each experiment are given in Table 4. A ReLU activation is used at the output of each neuron, except for the final one. Furthermore, the margin for the triplet loss is $\alpha=1$.
Figure 5: A visual depiction of the architecture we use for each sub-network $f$. The input layer has the same number of neurons as the dimensionality of the input. We use two hidden layers with, followed by an output layer with a single neuron. This final neuron is the latent space neuron that we interpret in our examples.
Table 4: Training Hyperparameters
Exp. No. Learning Rate Weight Decay Batch Size Input Size Hidden Size Output Size Epochs Factor Patience
1 0.001 0.0001 256 4 256 1 100 0.2 10
2 0.001 0.0001 256 4 256 1 300 0.2 10
3 0.001 0.0001 256 9 256 1 100 0.2 10
4 0.001 0.0001 256 9 256 1 100 0.2 10
5 0.001 0.0001 256 16 256 1 300 0.2 10
6 0.001 0.0001 256 6 256 1 300 0.5 10
7 0.001 0.0001 256 2 256 1 300 0.5 10
8 0.001 0.0001 256 2 256 1 300 0.5 10
9 0.001 0.0001 256 2 256 1 300 0.5 10
10 0.001 0.0001 256 2 256 1 300 0.5 10
11 0.001 0.0001 256 4 256 1 300 0.5 10
12 0.001 0.0001 256 4 256 1 300 0.5 10
Appendix E Dataset Generation
We retrieve the invariants of matrices and various physical systems using our method. We consider invariants of matrices under similarity and Lorentz transformations. Additionally, we investigate dynamical systems characterized by a variety of potentials, as well as the invariants in Minkowski spacetime. We choose the mass $m=1$ and spring constant $k=1$ where applicable.
E.1 Datasets and Training
E.1.1 Invariants Under the Similarity Transformation
In experiments 1-3, 5 in Table 1, we search for the trace and determinant of matrices under the similarity transformation. Each data point is a triplet consisting of three matrices of dimension $n$: an anchor matrix $A$, a positive example $P$, and a negative example $N$. The anchor is sampled by generating a random matrix. Each entry is sampled from a uniform distribution between $\left[\alpha,\beta\right]$. We try $\left[0,1\right]$, and $\left[-4,4\right]$. Neither choice affects the model’s ability to learn the invariant.
E.1.2 Invariants of Antisymmetric Matrices
For antisymmetric matrices in experiment 4, we prepare our dataset in the same way as we describe in E.1.1. We first sample an antisymmetric $3\times 3$ matrix for the anchor $A$, followed by a similarity transformation for the positive sample $P$. Finally, we sample a new antisymmetric matrix for the negative sample $N$. While both the anchor and negative samples are antisymmetric, the positive sample does not inherit this property under the transformation $P=MAM^{-1}$ when $M$ is not orthonormal, because antisymmetry is not preserved under a general change of basis. Hence, we use all 9 entries of the matrix as input, although we acknowledge that one could easily enforce that $M$ is orthonormal, in which case only 3 inputs would be needed from each of $A$, $P$, and $N$.
Since we use the antisymmetric anchor matrix $A$ as input when computing $\nabla_{\mathbf{x}}f(\mathbf{x})$, we expect that the result of symbolic search would simplify to the invariant in C.1.2, which is invariant under the similarity transformation.
The entries of each matrix are sampled from a normal distribution with $\mu=0$ and $\sigma=1$.
E.1.3 Invariants Under the Lorentz Transformation
In experiment 6, we apply the Lorentz transformation to the field strength tensor $F_{\mu\nu}$, which gives rise to the Lorentz invariants in C.1.3. Since the antisymmetry of $F_{\mu\nu}$ is preserved under the Lorentz transformation, each member of a triplet is antisymmetric, so we only use the 6 off-diagonal entries above (or equivalently below) the main diagonal as our input to the neural network. The anchor is a vector of these 6 entries from $F_{\mu\nu}$.
The entries of each matrix are sampled from a uniform distribution between $\left[0,1\right]$.
E.1.4 Potentials
The experiments in Table 2 correspond to motion in a potential, where we simulate trajectories by randomly sampling initial positions and velocities, and subsequently evolve these systems according to Hamilton’s equations. For each triplet $(x_{A},x_{P},x_{N})$, the anchor $x_{A}$ and positive sample $x_{P}$ are measurements at two different points along the same trajectory, while the negative sample $x_{N}$ is sampled from a different trajectory. The network must determine whether or not two measurements belong to the same particle. See C.3 for details regarding the invariants.
The dataset is generated with mass $m=1$, spring constant $k=1$, and a time grid $t\in[0,5]$ with 10,001 points. Initial conditions are sampled from $0,1]^{2}$, and trajectory points are selected using random indices $i,j$ from the solution.
E.1.5 Spacetime
In experiment 12 in Table 3, each triplet again consists of an anchor $x_{A}$, a positive sample $x_{P}$, and a negative sample $x_{N}$. The anchor is a randomly sampled four-vector representing an event in Minkowski spacetime. Each entry is sampled from a uniform distribution between $\left[0,1\right]$. The positive sample is generated by applying a Lorentz transformation to the anchor, ensuring that the spacetime interval remains invariant. The negative sample, on the other hand, is another randomly generated four-vector that does not share the same spacetime interval as the anchor, allowing the neural network to distinguish between vectors that do and do not preserve this invariant.
Appendix F Results of Direct Symbolic Regression
We compare our method to direct symbolic regression, which is can only recover the ground truth if the latent space encodes it in a linear manner. In this case, we obtain latents from the trained neural network, and supply them to the standard symbolic regression algorithm Cranmer (2023), which then attempts to extract a symbolic equation from the dataset consisting of the pairs $(X,\text{latents})$. In Table 5, we apply direct symbolic regression on the same experiments as in tables 1-3. We report the retrieved expression only if the correct expression is retrieved. Notably, only 7 out of 12 experiments succeed with direct symbolic regression. We also note that in the symbolic gradient experiments, we report the first result. In symbolic regression we allowed for multiple runs, including up to $200000$ data points. It is interesting to see that in one case, symbolic regression could recover a ground truth equation while failing to approximate the latents Figure 2.
Table 5: Results of Direct Symbolic Regression
Exp. No. Expression Retrieved? Expression
1 Yes $\frac{A_{11}+A_{22}}{-0.12}$
2 Yes $A_{21}A_{12}-A_{22}A_{11}$
3 Yes $A_{11}+A_{22}+A_{33}-1.489$
4 Yes $3.239-A_{12}A_{12}-A_{23}A_{23}-A_{13}A_{13}$
5 Yes $A_{11}+A_{22}+A_{33}+A_{44}-1.907$
6 No -
7 Yes $394.111x^{2}-(-399.431)v^{2}$
8 No -
9 No -
10 No -
11 Yes $x_{2}v_{1}-x_{1}v_{2}$
12 No -
Figure 6: Direct symbolic regression on the latents typically fails when the concept is encoded in a non-linear manner. However, in the peculiar case of experiment 4 in Table 5, this method retrieves the ground truth expression. This is likely because it is fitting to a data-dense region of the non-linear plot, as shown in this figure. The blue line represents the equation extracted through direct symbolic regression. It almost perfectly passes through a linear sub-region of the correlation curve.